# A Miyaoka-Yau inequality for hyperplane arrangements in $\mathbb{CP}^{n}$

### v1

@@ 3.3 Polarization
[…]
@@
where $f=\sum_{j=0}^{n-2}C_{j}k^{j}$ is a polynomial in $k$ of degree at most $n-2$.
@@
with $\mathcal{S}$ a set of not necessarily distinct $n-j$ irreducible subspaces $L^{\prime}\in\mathcal{L}_{\textup{irr}}$, which only depends on $\mathcal{H}$. ∎
@@
where $\omega_{\operatorname{FS}}$ is the Fubini-Study metric normalized so that $[\omega_{\operatorname{FS}}]\in c_{1}(\mathcal{O}_{\mathbb{P}^{n}}(1))$ and $\eta$ is a fixed real and closed $(1,1)$-form. The form $\eta$ depends only on the polarization coefficients and it is positive when restricted to the kernel of $\pi_{*}$, it can be constructed as in the proof of Proposition 3.24 in [Voi07a].
@@ 4.1 General theory
Let $X$ be a complex manifold and let $D$ be a simple normal crossing divisor with irreducible decomposition $D=\bigcup_{i\in I}D_{i}$ Let $\mathcal{E}$ be a holomorphic vector bundle on $X$.
@@
    $F^{i}_{a}=\mathcal{E}^{i}_{a}|_{D_{i}}\,\,\big{/}\,\,\mathcal{E}(-D_{i})|_{D_{i}}$ (4.1)
@@
Since the sheaves $\mathcal{E}^{i}_{a}$ and $\mathcal{E}(-D_{i})$ are locally free $\mathcal{O}_{X}$-modules, the restrictions $\mathcal{E}^{i}_{a}|_{D_{i}}$ and $\mathcal{E}(-D_{i})|_{D_{i}}$ are locally free $\mathcal{O}_{D_{i}}$-modules. The quotient (4.1) is a subsheaf of
$\mathcal{E}|_{D_{i}}=\mathcal{E}|_{D_{i}}\,\,\big{/}\,\,\mathcal{E}(-D_{i})|_{D_{i}}.$
Item (iii) requires that the $F^{i}_{a}$ are locally free $\mathcal{O}_{D_{i}}$-modules and that for $a^{\prime}<a$ the inclusions of $\mathcal{O}_{D_{i}}$-modules
$F^{i}_{a^{\prime}}\subset F^{i}_{a}\subset\mathcal{E}|_{D_{i}}$
are inclusions of vector subbundles.
@@
where $\operatorname{rank}(\operatorname{Gr}^{i}_{a})$ is the rank of the vector bundle $\operatorname{Gr}^{i}_{a}$ on $D_{i}$
@@
    For any $\mathbf{a}=(a_{1},\ldots,a_{k})\in\mathbb{R}^{k}$, the intersection $F_{\mathbf{a}}$ is a vector bundle on $D_{J}$ ; and the sum $\sum_{\mathbf{a}^{\prime}\lneq\mathbf{a}}F_{\mathbf{a}^{\prime}}$ is a vector subbundle of $F_{\mathbf{a}}$ [^5] For generic values of $\mathbf{a}$, e.g. if $\mathbf{a}$ does not belong to the finite set $\prod_{j=1}^{k}\operatorname{wt}(i_{j})$, the subbundle $\sum_{\mathbf{a}^{\prime}\lneq\mathbf{a}}F_{\mathbf{a}^{\prime}}$ is actually equal to $F_{\mathbf{a}}$.
@@
If $\mathcal{E}_{*}$ is locally abelian, then we can decompose $\mathcal{E}=\oplus_{\mathbf{a}}U_{\mathbf{a}}$ locally on $D_{J}$ as in Definition 4.8 . Choose a frame $S_{\mathbf{a}}$ for each component $U_{\mathbf{a}}$ and let $S=\cup_{\mathbf{a}}S_{\mathbf{a}}$. It is easy to see that $S$ is adapted to $\mathcal{E}_{*}$
@@
Finally, $\operatorname{rank}(\operatorname{Gr}^{i,j}_{a,b})$ denotes the rank of the vector bundle $\operatorname{Gr}^{i,j}_{a,b}$
@@ 4.2 Weights and filtration
[…]
@@
Note that $10<a_{L}<$ 1 because of the klt condition (1.1).
@@ 4.5 The parabolic second Chern character
$\nu_{L,M}=\begin{cases}\operatorname{rank}\big{(}\operatorname{Gr}^{L,M}_{a_{L},a_{M}}\big{)}&\textup{ if }L\neq M\,,\\ \operatorname{rank}\big{(}\operatorname{Gr}^{L}_{a_{L}}\big{)}=\operatorname{codim}L&\textup{ if }L=M\,,\end{cases}$
@@
where $\alpha=r\cdot\pi^{*}h\in H^{2}(X,\mathbb{Z})$ and $\imath^{*}$ is the pullback by the inclusion map of $D_{L}\subset X$. The projection formula $\imath_{*}\left(\imath^{*}\alpha\right)=\alpha\cdot\gamma_{L}$see Equation 1.6 in [Voi07b]- implies that
@@
Let $L$ and $M$ be irreducible subspaces with $D_{L}\cap D_{M}\neq\emptyset$ Then
@@
$\begin{gathered}\operatorname{par-ch}_{2}(\mathcal{E}_{*})=-\frac{n+1}{2}\cdot(\pi^{*}h)^{2}\\ +\,\frac{1}{2}\sum_{L\in\mathcal{L}_{\textup{irr}}}a_{L}^{2}\cdot\operatorname{codim}L\cdot\gamma_{L}^{2}+\sum_{L\subsetneq M}a_{L}a_{M}\cdot\operatorname{codim}M\cdot\gamma_{L}\cdot\gamma_{M}\,,\end{gathered}$ (4.33)
@@ 5 Stability of the parabolic bundle
where $\mathcal{V}_{*}$ is the parabolic structure on $\mathcal{V}$ induced from $\mathcal{E}_{*}$
@@ 5.1 Auxiliary results
Let $\mathcal{V}\subset T\mathbb{CP}^{n}$ be a distribution and let $H\subset\mathbb{CP}^{n}$ be a hyperplane. We say that $\mathcal{V}$ is tangent to $H$ (or that $H$ is tangent to $\mathcal{V}$), if for every point $x\in H\cap U$where $U$ is the regular set of $\mathcal{V}$- we have
@@
Let $M\subset\mathbb{CP}^{n}$ be a linear subspace with $\dim M=r-1$ for some $11\leq r\leq n-$ 1. The collection of all $r$-dimensional subspaces that contain $M$ defines a distribution $\mathcal{V}\cong\mathcal{O}_{\mathbb{P}^{n}}(1)^{\oplus r}$ of index $\imath=r$. A hyperplane $H$ belongs to $\operatorname{Tan}(\mathcal{V})$ if and only if $H\supset M$. In particular,
@@
has homogeneous polynomial coefficients of degree $3$ and the contraction of $\omega$ with the Euler vector field on $\mathbb{C}^{4}$ is identically zero. The $1$-form $\omega$ defines a codimension $1$ distribution $\mathcal{V}$ on $\mathbb{CP}^{3}$ of rank $r=2$ and degree $d=2$ (see Remark A.69). The index of $\mathcal{V}$ is equal to $\imath=r-d=0$. The distribution $\mathcal{V}$ is integrable. The leaves of the foliation defined by $\mathcal{V}$ make a pencil of cubic surfaces in $\mathbb{CP}^{3}$ given by $\{\lambda_{1}\cdot f+\lambda_{2}\cdot h^{3}=0\}$ with $[\lambda_{1},\lambda_{2}]\in\mathbb{CP}^{1}$. A holomorphic vector field $Y$ tangent to $\mathcal{V}$ must also be tangent to the smooth cubic surface $\{f=0\}$ and therefore it must vanish.
@@ 5.2 Key estimate
The number $\delta>0$ in Proposition 5.10 depends only on the weighted arrangement $(\mathcal{H},\mathbf{a})$as given by Equation (5.11)- and not on the distribution $\mathcal{V}$.
@@
We denote by $O(k^{n-2})$ a polynomial in $k$ of degree at most $n-2$, say $\sum_{j=0}^{n-2}C_{j}k^{j}$, such that there is a positive number $K=K(\mathcal{H},b_{L})$, depending only on $\mathcal{H}$ and the polarization coefficients $b_{L}$, such that $|C_{j}|\leq K$ for all $j$.
@@
$\begin{gathered}\operatorname{par-deg}_{P_{k}}(\mathcal{V}_{*})=\deg_{P_{k}}(\mathcal{V})\,\,-\\ \left(\sum_{H|H\pitchfork\mathcal{V}}a_{H}\cdot c_{1}(P_{k})^{n-1}\cdot\gamma_{H}+\sum_{L\in\mathcal{L}_{\textup{irr}}^{\circ}}r_{L}a_{L}\cdot c_{1}(P_{k})^{n-1}\cdot\gamma_{L}\right)\,,\end{gathered}$ (5.15)
@@
is $<0$. Since index of the distribution $\mathcal{V}\subset T\mathbb{CP}^{n}$ is equal to $\imath\geq 0$, by Proposition 5.10 we get that $c=-\delta$ for some $\delta>0$ that depends only on $(\mathcal{H},\mathbf{a})$. This concludes the proof of Theorem 5.1 . ∎
@@
The expression $p(k)$ defines a polynomial in $k$ of degree $n-2$ with real coefficients. The coefficients of $p(k)$ depend only on the weighted arrangement $(\mathcal{H},\mathbf{a})$ and the fixed integers $b_{L}$ involved in the choice of polarization $P_{k}$. More precisely, if we write
@@ 6.1 The quadratic form of an arrangement
$Q(\mathbf{a})=4(n+1)\cdot\sum_{L\in\mathcal{L}_{\textup{irr}}^{n-2}}a_{L}^{2}\,-\,2(n+1)\cdot\sum_{H\in\mathcal{H}}B_{H}\cdot a_{H}^{2}\,-\,2\cdot s^{2}\,,$ (6.3)
@@
$Q_{ij}=\begin{cases}-(n+1)\sigma_{i}+2n&\text{ if }i=j\,,\\ -2&\text{ if }i\neq j\text{ and }H_{i}\cap H_{j}\text{ is reducible },\\ \,\,n-1&\text{ if }i\neq j\text{ and }H_{i}\cap H_{j}\text{ is irreducible }\,.\end{cases}$ (6.5)
@@
Let =$\mathcal{H}=\{H_{1},\ldots,H_{N}\}$ be an arrangement in $\mathbb{CP}^{n}$ which is generic, in the sense that no $3$ distinct hyperplanes are linearly dependent. Suppose that the number of hyperplanes $N$ is $>n+1$. The generic assumption implies that $\sigma_{i}=0$ for all $i$. Equation (6.5) gives us
@@
The assumption that $N>n+1$ implies that the eigenvalue is $<0$. On the other hand, if $\mathbf{a}=(a_{1},\ldots,a_{N})$ is a vector with $\sum_{i}a_{i}=0$, i.e., $\mathbf{a}$ is orthogonal to $\mathbf{1}$, then
@@
This shows that the orthogonal complement $\mathbf{1}^{\perp}$ is an eigenspace of $Q$ with positive eigenvalue. Thus, the quadratic form $Q$ is non-degenerate of signature $(N-1,1)$.
@@
Let $\mathcal{H}\subset\mathbb{CP}^{n}$ be the braid (or $A_{n+1}$) arrangement consisting of the $\binom{n+2}{2}$ hyperplanes $H_{ij}=\{x_{i}=x_{j}\}$ for $21\leq i<j\leq n+$ 2 in
@@
$(\mathbf{q}_{ij})_{kl}=\begin{cases}n(1-n)&\text{ if }\{k,l\}=\{i,j\}\,,\\ -2&\text{ if }\{k,l\}\cap\{i,j\}=\emptyset\,,\\ n-1&\text{ if }|\{k,l\}\cap\{i,j\}|=1\,.\end{cases}$ (6.9)
@@
$(\mathbf{v}_{i})_{kl}=\begin{cases}1\,\,&\textup{ if }\,\,i\in\{k,l\}\\ 0\,\,&\textup{ if }\,\,i\notin\{k,l\}\end{cases}$
@@ 6.2 The stable cone
Let $\mathcal{H}\subset\mathbb{CP}^{n}$ be the braid arrangement as in Example 6.7. The irreducible subspaces $L_{I}$ correspond to subsets $I\subset[n+2]$ with $12\leq|I|\leq n+$ 1 by letting $L_{I}=\{x_{i}=x_{j}\,\,\textup{ for }\,\,i,j\in I\}$. In particular, $\operatorname{codim}L_{I}=|I|-1$. By Lemma 6.14 (i), the weighted arrangement $(\mathcal{H},\mathbf{a})$ with weights $a_{ij}>0$ is stable if and only if for every $I\subset[n+2]$ with $12\leq|I|\leq n+$ 1 we have
@@ 6.3 The matroid polytope and the semistable cone
$(\mathbf{e}_{\mathcal{B}})_{H}=\begin{cases}1&\text{ if }H\in\mathcal{B}\,,\\ 0&\text{ if }H\notin\mathcal{B}\,.\end{cases}$
@@
where $k$ is the number of factors in the decomposition of $\mathcal{H}$ as a product of irreducible arrangements $\mathcal{H}\cong\mathcal{H}_{1}\times\ldots\times\mathcal{H}_{k}$

### v2

@@ 3.3 Polarization
Let $X\xrightarrow{\pi}\mathbb{CP}^{n}$ be the resolution of $\mathcal{H}$ as in Definition 3.1 . Recall that we write
@@
where $f=\sum_{j=0}^{n-2}C_{j}k^{j}$ is a polynomial in $k$ of degree at most n2n-2.
@@
with $\mathcal{S}$ a set of not necessarily distinct njn-j irreducible subspaces $L^{\prime}\in\mathcal{L}_{\textup{irr}}$, which only depends on $\mathcal{H}$. ∎
@@
where $\omega_{\operatorname{FS}}$ is the Fubini-Study metric normalized so that $\omega_{\operatorname{FS}}]\in c_{1}(\mathcal{O}_{\mathbb{P}^{n}}(1))$ and $\eta$ is a fixed real and closed $(1,1)$-form. The form $\eta$ depends only on the polarization coefficients and it is positive when restricted to the kernel of $\pi_{*}$, it can be constructed as in the proof of Proposition 3.24 in [Voi07a].
@@ 4.1 General theory
Let $X$ be a complex manifold and let $D$ be a simple normal crossing divisor with irreducible decomposition $D=\bigcup_{i\in I}D_{i}$ . Let $\mathcal{E}$ be a holomorphic vector bundle on $X$.
@@
    $F^{i}_{a}=\left(\mathcal{E}^{i}_{a}\,\,\big/\,\,\mathcal{E}(-D_{i})\right)|_{D_{i}}$ (4.1)
@@
Item (iii) requires that the sheaf $F^{i}_{a}$ obtained by restriction of the skyscraper sheaf $\mathcal{E}^{i}_{a}\,\,\big/\,\,\mathcal{E}(-D_{i})$ to $D_{i}$ is a locally free $\mathcal{O}_{D_{i}}$-module. Moreover, the inclusion of $\mathcal{O}_{D_{i}}$-modules
$F^{i}_{a}\subset\mathcal{E}|_{D_{i}}=\left(\mathcal{E}\,\,\big/\,\,\mathcal{E}(-D_{i})\right)|_{D_{i}}$
is an inclusion of vector bundles.
@@
where $\operatorname{rank}(\operatorname{Gr}^{i}_{a})$ is the rank of the vector bundle $\operatorname{Gr}^{i}_{a}$ on $D_{i}$ .
@@
    For any $\mathbf{a}=(a_{1},\ldots,a_{k})\in\mathbb{R}^{k}$, the intersection $F_{\mathbf{a}}$ is a vector bundle on $D_{J}$ ; and the sum $\sum_{\mathbf{a}^{\prime}\lneq\mathbf{a}}F_{\mathbf{a}^{\prime}}$ is a vector subbundle of $F_{\mathbf{a}}$ . [^5] For generic values of $\mathbf{a}$, e.g. if $\mathbf{a}$ does not belong to the finite set $\prod_{j=1}^{k}\operatorname{wt}(i_{j})$, the subbundle $\sum_{\mathbf{a}^{\prime}\lneq\mathbf{a}}F_{\mathbf{a}^{\prime}}$ is actually equal to $F_{\mathbf{a}}$.
@@
If $\mathcal{E}_{*}$ is locally abelian, then we can decompose $\mathcal{E}=\oplus_{\mathbf{a}}U_{\mathbf{a}}$ locally on $D_{J}$ as in Definition 4.8 . Choose a frame $S_{\mathbf{a}}$ for each component $U_{\mathbf{a}}$ and let $S=\cup_{\mathbf{a}}S_{\mathbf{a}}$. It is easy to see that $S$ is adapted to $\mathcal{E}_{*}$ .
@@
Finally, $\operatorname{rank}(\operatorname{Gr}^{i,j}_{a,b})$ denotes the rank of the vector bundle $\operatorname{Gr}^{i,j}_{a,b}$ .
@@ 4.2 Weights and filtration
Let $(\mathcal{H},\mathbf{a})$ be a weighted arrangement as in Theorem 1.1 . Let $X\xrightarrow{\pi}\mathbb{CP}^{n}$ be the resolution of $\mathcal{H}$ and let $\mathcal{E}$ be the pullback tangent bundle
@@
Note that $0<a_{L}<1$ because of the klt condition (1.1).
@@ 4.5 The parabolic second Chern character
[…]
@@
where $\alpha=r\cdot\pi^{*}h\in H^{2}(X,\mathbb{Z})$ and $\imath^{*}$ is the pullback by the inclusion map of $D_{L}\subset X$. The projection formula $\imath_{*}\left(\imath^{*}\alpha\right)=\alpha\cdot\gamma_{L}$ -see Equation 1.6 in [Voi07b]- implies that
@@
Let $L$ and $M$ be irreducible subspaces with $D_{L}\cap D_{M}\neq\emptyset$ . Then
@@
[…]
@@ 5 Stability of the parabolic bundle
where $\mathcal{V}_{*}$ is the parabolic structure on $\mathcal{V}$ induced from $\mathcal{E}_{*}$ .
@@ 5.1 Auxiliary results
Let $\mathcal{V}\subset T\mathbb{CP}^{n}$ be a distribution and let $H\subset\mathbb{CP}^{n}$ be a hyperplane. We say that $\mathcal{V}$ is tangent to $H$ (or that $H$ is tangent to $\mathcal{V}$), if for every point $x\in H\cap U$ -where $U$ is the regular set of $\mathcal{V}$- we have
@@
Let $M\subset\mathbb{CP}^{n}$ be a linear subspace with $\dim M=r-1$ for some $1\leq r\leq n-1$. The collection of all $r$-dimensional subspaces that contain $M$ defines a distribution $\mathcal{V}\cong\mathcal{O}_{\mathbb{P}^{n}}(1)^{\oplus r}$ of index $\imath=r$. A hyperplane $H$ belongs to $\operatorname{Tan}(\mathcal{V})$ if and only if $H\supset M$. In particular,
@@
has homogeneous polynomial coefficients of degree $3$ and the contraction of $\omega$ with the Euler vector field on $\mathbb{C}^{4}$ is identically zero. The $1$-form $\omega$ defines a codimension $1$ distribution $\mathcal{V}$ on $\mathbb{CP}^{3}$ of rank $r=2$ and degree $d=2$ (see Remark A.69). The index of $\mathcal{V}$ is equal to $\imath=r-d=0$. The distribution $\mathcal{V}$ is integrable. The leaves of the foliation defined by $\mathcal{V}$ make a pencil of cubic surfaces in $\mathbb{CP}^{3}$ given by $\{\lambda_{1}\cdot f+\lambda_{2}\cdot h^{3}=0\}$ with $\lambda_{1},\lambda_{2}]\in\mathbb{CP}^{1}$. A holomorphic vector field $Y$ tangent to $\mathcal{V}$ must also be tangent to the smooth cubic surface $\{f=0\}$ and therefore it must vanish.
@@ 5.2 Key estimate
The number $\delta>0$ in Proposition 5.10 depends only on the weighted arrangement $(\mathcal{H},\mathbf{a})$ -as given by Equation (5.11)- and not on the distribution $\mathcal{V}$.
@@
We denote by $O(k^{n-2})$ a polynomial in $k$ of degree at most n2n-2, say $\sum_{j=0}^{n-2}C_{j}k^{j}$, such that there is a positive number $K=K(\mathcal{H},b_{L})$, depending only on $\mathcal{H}$ and the polarization coefficients $b_{L}$, such that $|C_{j}|\leq K$ for all $j$.
@@
[…]
@@
is <0<0. Since index of the distribution $\mathcal{V}\subset T\mathbb{CP}^{n}$ is equal to $\imath\geq 0$, by Proposition 5.10 we get that $c=-\delta$ for some $\delta>0$ that depends only on $(\mathcal{H},\mathbf{a})$. This concludes the proof of Theorem 5.1 . ∎
@@
The expression $p(k)$ defines a polynomial in $k$ of degree n2n-2 with real coefficients. The coefficients of $p(k)$ depend only on the weighted arrangement $(\mathcal{H},\mathbf{a})$ and the fixed integers $b_{L}$ involved in the choice of polarization $P_{k}$. More precisely, if we write
@@ 6.1 The quadratic form of an arrangement
[…]
@@
[…]
@@
Let $\mathcal{H}=\{H_{1},\ldots,H_{N}\}$ be an arrangement in $\mathbb{CP}^{n}$ which is generic, in the sense that no $3$ distinct hyperplanes are linearly dependent. Suppose that the number of hyperplanes $N$ is >n+1>n+1. The generic assumption implies that $\sigma_{i}=0$ for all $i$. Equation (6.5) gives us
@@
The assumption that $N>n+1$ implies that the eigenvalue is <0<0. On the other hand, if $\mathbf{a}=(a_{1},\ldots,a_{N})$ is a vector with $\sum_{i}a_{i}=0$, i.e., $\mathbf{a}$ is orthogonal to $\mathbf{1}$, then
@@
This shows that the orthogonal complement $\mathbf{1}^{\perp}$ is an eigenspace of $Q$ with positive eigenvalue. Thus, the quadratic form $Q$ is non-degenerate of signature (N1,1)(N-1,1).
@@
Let $\mathcal{H}\subset\mathbb{CP}^{n}$ be the braid (or $A_{n+1}$) arrangement consisting of the $\binom{n+2}{2}$ hyperplanes $H_{ij}=\{x_{i}=x_{j}\}$ for $1\leq i<j\leq n+2$ in
@@
where $\langle\cdot,\cdot\rangle$ is the Euclidean product on $\mathbb{R}^{N}$ and the vectors $\mathbf{q}_{ij}$ are the rows of the matrix $Q$ given by Equation (6.5)
@@
[…]
@@ 6.2 The stable cone
Let $\mathcal{H}\subset\mathbb{CP}^{n}$ be the braid arrangement as in Example 6.7. The irreducible subspaces $L_{I}$ correspond to subsets $I\subset[n+2]$ with $2\leq|I|\leq n+1$ by letting $L_{I}=\{x_{i}=x_{j}\,\,\textup{ for }\,\,i,j\in I\}$. In particular, $\operatorname{codim}L_{I}=|I|-1$. By Lemma 6.14 (i), the weighted arrangement $(\mathcal{H},\mathbf{a})$ with weights $a_{ij}>0$ is stable if and only if for every $I\subset[n+2]$ with $2\leq|I|\leq n+1$ we have
@@ 6.3 The matroid polytope and the semistable cone
[…]
@@
where $k$ is the number of factors in the decomposition of $\mathcal{H}$ as a product of irreducible arrangements $\mathcal{H}\cong\mathcal{H}_{1}\times\ldots\times\mathcal{H}_{k}$ .